OpenStudy (anonymous):

solve by substitution 9xt5y=-4 x=34-5y

OpenStudy (dumbcow):

9x +5y = -4 x = 34-5y ??

OpenStudy (anonymous):

yes thats it

OpenStudy (dumbcow):

ok since x is already solved for in bottom equation, substitute it into top equation \[9(34-5y) +5y = -4\] \[306-45y +5y = -4\] \[-40y = -310\] \[y = 310/40 = 31/4\] then plug that into bottom equation to solve for x

OpenStudy (anonymous):

it is an ordered pair correct?

OpenStudy (dumbcow):

yes, solution will be in the form (x,y)

OpenStudy (anonymous):

x okay I give got an answer because all I get are -4, 13

OpenStudy (dumbcow):

give it another try \[x = 34 - 5(\frac{31}{4})\]

OpenStudy (anonymous):

29,7

OpenStudy (anonymous):

sorry I am just dumb at this

OpenStudy (dumbcow):

where are you getting these numbers from? well where are you struggling....break it down into simple steps whats 5*(31/4) ?

OpenStudy (anonymous):

31/20

OpenStudy (dumbcow):

think of the 5 as (5/1) when you multiply fractions, multiply straight across, top times top, bottom times bottom 5*(31/4) = (5*31)/(1*4) make sense?

OpenStudy (anonymous):

first=155 second 4

OpenStudy (dumbcow):

good, next part subtracting fractions, common denominator of 4 34 - 155/4

OpenStudy (anonymous):

I thought you did them top to bottom no wonder I am failing

OpenStudy (anonymous):

30 151

OpenStudy (dumbcow):

no never do top to bottom when multiplying....you prob get confused with cross-multiplying when there is an equal sign 5/2 = 4/5

OpenStudy (dumbcow):

ok looks like you subtracted 4 from each number?

OpenStudy (anonymous):

Never did any of this in school did not have to but yes I subt that way

OpenStudy (dumbcow):

so you don't know how to add/subtract fractions

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

sorry you picked the dummy to help

OpenStudy (dumbcow):

no thats fine but you are going to need a lot of help, since it looks like you are in an algebra class....you can't do algebra before you know how to do fractions

OpenStudy (anonymous):

thanks I tried to tell them that but, as you can see it did not work! I am doing online and am 50 yrs old so yeah I feel real stupid

OpenStudy (dumbcow):

oh well there are a lot of online resources and books on basic math so you can teach yourself and get caught up. With fractions, just remember that the bottom number needs to be the same before you can add/subtract \[\frac{34}{1} - \frac{155}{4}\] change the 1st fraction so it has a 4 on bottom, do that by multiplying by (4/4) \[\frac{34}{1}*\frac{4}{4} - \frac{155}{4}\] \[\frac{136}{4} - \frac{155}{4}\] now just subtract the top numbers and leave bottom alone \[\frac{136}{4} -\frac{155}{4} = \frac{-19}{4}\]

OpenStudy (dumbcow):

solution: (-19/4, 31/4)

OpenStudy (dumbcow):

here is a website that might help you http://www.mathexpression.com/learning-fractions.html